# Building Machine Learning Challenges for Anomaly Detection in Science

### v1

@@ Abstract
Scientific discoveries are often made by finding a pattern or object that was not predicted by the known rules of science. Oftentimes, these anomalous events or objects that do not conform to the norms are an indication that the rules of science governing the data are incomplete, and something new needs to be present to explain these unexpected outliers. The challenge of finding anomalies can be confounding since it requires codifying a complete knowledge of the known scientific behaviors and then projecting these known behaviors on the data to look for deviations. When utilizing machine learning, this presents a particular challenge since we require that the model not only understands scientific data perfectly but also recognizes when the data is inconsistent and out of the scope of its trained behavior. In this paper, we present three datasets aimed at developing machine learning-based anomaly detection for disparate scientific domains covering astrophysics, genomics, and polar science. We present the different datasets along with a scheme to make machine learning challenges around the three datasets findable, accessible, interoperable, and reusable (FAIR). Furthermore, we present an approach that generalizes to future machine learning challenges, enabling the possibility of large, more compute-intensive challenges that can ultimately lead to scientific discovery.
@@ 1 Introduction
Anomaly detection, the process of finding an outlier in a distribution, has long been a critical element of scientific discovery. New anomalies often trigger the question of how to construct a model that includes everything allowing for a deeper understanding of the data and, ultimately, a better scientific model. In some cases, this has happened with the observation of a new event or process. Other times, discoveries have been made through aggregating data to observe unexplainable trends [1, 2, 3, 4].
@@
In this paper, we present a series of machine learning (ML) challenges targeting anomaly detection in different scientific domains. We define this as the ability to identify events, objects, or phenomena that are rare and do not match ordinary detector noise or otherwise known signals or phenomena. Anomaly detection is thus searching for the “unknown,” including “known-unknowns” (i.e., things that may have been hypothesized and anticipated but otherwise poorly modeled or quantified) and “unknown-unknowns” (i.e., things that can not be accommodated by any prior knowledge in any given domain and which—if detected—will push the frontier in understanding the theory and phenomenology of the underlying observations and measurements). This could be the identification of poorly modeled phenomena, such as a core-collapse supernova signal in a gravitational-wave detector[5, 6, 7, 8, 9], unexplained sea ice retreat [10], extreme snow melts in the Arctic [11], or the identification of an unexplained object, such as the observation of a new astrophysical transient[12], a new type of material or a new species of animal [13, 14, 15].The goal of the challenges presented here is to bring awareness to the difficulty of using ML for these types of problems: the options of well-labeled datasets or even the possibility of training through a systematic procedure, such as token masking, are often insufficient to answer this type of question. Despite that, there is tremendous potential for algorithms focused on this strategy to enable scientific discovery.
@@
The process of using ML for anomaly detection, sometimes referred to by the less inclusive term outlier detection, has existed for a significant amount time [16, 17, 18, 19]. However, the utilization of anomaly detection within Artificial Intelligence (AI) has been difficult due to the ambiguity in labeling; this has been particularly true in the creation of training strategies and loss functions that indicate both the level to which an object is anomalous and the confidence that this anomaly is a scientifically meaningful deviation[20, 21]. Nevertheless, a number of approaches have been tried  [21, 19, 12], including the use of unsupervised algorithms [22], which aim to build a mapping for what is regular through an embedding in a latent space, semi-supervised algorithms, which aim to use partially labeled data to construct representation space characteristic to the problem at hand [23], and weakly-supervised algorithms, which aim to play regions of data off of each other to find the largest incongruous regions of the data[24].
@@
This paper, the ML challenge project, aims to unite disparate scientific communities from many different domains under the umbrella of a single, challenging scientific problem. The communities involved are part of the NSF Harnessing the Data Revolution (HDR) program, which aims to connect domain science with AI to enable groundbreaking scientific discoveries. In the interest of connecting most of the HDR communities, we engaged all five HDR institutes to develop datasets with a tractable problem that we define across the different domains. For each institute, we target a single anomaly detection problem and dataset. Furthermore, we aim to ensure that the anomaly detection problems are sufficiently aligned such that the same algorithmic toolkit can be applied to each of the challenges. As a result of this, we narrowed the challenge from five initially proposed datasets to three. Lastly, the most critical component of this challenge was the ability to hide pre-identified anomalies in such a way that there was no ambiguity in the anomaly definition in the final dataset. The challenge is to see if the proponents can find these pre-hidden anomalies with the smallest possible false positive rate. It is the goal of each team to figure out either a single algorithm or series of algorithms that can be applied to each of the datasets to find the anomalies. The three challenges cover a wide variety of datasets and problems ranging from a challenge on time series data, to one on photographic images, and a third on satellite imagery over time.
A diagram of the challenge flow is shown in figure 1. To perform the challenge, teams must develop an algorithm that reads the data and provides a score for the likelihood that qualifies a data point is an anomaly. The scoring is then determined by seeing if the algorithm is capable of singling out the anomalous features of the dataset. To facilitate the challenge, we provide three core data subsets:
  • Training Dataset: This is the publicly released subset consisting of the background (aka non-anomalous) cases and possible hypothetical anomalous signals to enable the model construction. The hypothetical signals do not need to include the morphologies of the challenge dataset. They just act as possible examples of what might be present. Moreover, the background dataset need not be exactly the same proportional composition as the final dataset.
  • Testing Dataset: This subset is used for benchmarking the algorithm and presenting a leaderboard on the competition platform. It largely consists of background non-anomalous events with a few hidden anomalies. The anomalies are all known through additional information that is not available in the challenge. This dataset is never released publicly; however, participants may run their algorithm on it through the collaboration platform.
  • Challenge Dataset: This subset is used to determine the winner. It is constructed in the same way as the testing dataset. However, it is never released publicly nor is it used before the final scoring. The absence a score until the competition is decided ensures that no participating team can overly tune the algorithm on the dataset.
For the dataset creation, we adhered to standardized reproducibility, with the curation and data cleaning workflows stored on GitHub. The preparation of the private datasets will be released following the end of the challenge.
For the ML Challenge framework, we have chosen the Codabench platform [25]. Codabench is a flexible open-source platform designed for benchmarking ML algorithms, particularly through ML challenges, that enables custom code for scoring and presentation of the results. Codabench provides a front end with a standard interface to view scoring, the leaderboard, and instructions for how to participate in the challenge. The back end of the challenge uses the National Energy Research Scientific Computing Center (NERSC) at Lawrence Berkeley labs[26]. Submissions on the Codabench platform are forwarded to NERSC where the evaluation and scoring are performed.
Figure 1: Overview of the flow of the ml challenge, and the datasets used for training. All preparation was explicitly made under FAIR principles. The Codabench platform and NERSC computing cluster are used for scoring and ranking of submissions.
Findable, Accessible, Interoperable, and Reusable (FAIR) principles are fundamental to the ML challenge[27]. In the interest of ensuring reproducibility, members of the team have worked to ensure that the challenge adhered to full FAIR standards with dataset curation, appropriate metadata, and provenance. Additionally, we added the requirement that all submissions be constructed in a fully FAIR way, such that the submitted jobs were completely FAIR, and could be used on an arbitrary computing cluster. To ensure this, we added a requirement to the submissions to construct the model inference environment within the code scoring submission. Through the use of both NERSC and the insistence on a FAIR framework, we aim to democratize the ML Challenge through a common computing cluster that evaluates the scores, and environment construction that requires all used packages to be publicly accessible to compute nodes on NERSC. Further extensions can be made to ensure FAIR principles can be adapted to other aspects of the ML challenge, such as model construction, in future editions.
A driving goal of this work is to codify a list of best practices that ensure workflows, no matter the scientific domain, are easy to follow and reproduce. Through this work, we aim to establish an exemplar for future ML challenges, preserving the level of FAIRness present in this challenge in future challenges. Moreover, by ensuring a high level of FAIRness we aim to use these datasets for future benchmarking[20, 28, 29]. Established benchmarks can help drive the development of future algorithms, ultimately enhancing our ability to make scientific discoveries.
For this challenge, three of the five HDR institutes are utilized. These datasets consist of
@@
  • Anomalous Gravitational-Wave Transients: This consists of short-duration (O(1s)) astrophysical signals present in public data from the two LIGO gravitational-wave detectors. The data that is received from each of the LIGO interferometers is a gravitational-wave strain from a signal, indicating a fractional stretching and squeezing of space-time on the order of $10^{-22}$ the measured length scale. The challenge is to find a signal that appears in both detectors and has broad features representative of an astrophysical signature. Such a signature may correspond, for example, to a core-collapse supernova, a white noise burst of unknown origin, or a black hole merger beyond our nominal modeling and expectation from general relativity.
@@
  • Butterfly Hybrids: This dataset consists of images of butterflies from two mimetic species, each with highly varied subspecies [30]. It includes images of the various subspecies, as well as different “hybrids” among each species’ subspecies. The challenge is to find the anomalous butterfly hybrids without confusing them for subspecies of the other species.
@@
  • Anomalous Sea Level Rise: This dataset consists of satellite imagery [31] depicting observed sea levels over time. The challenge is to use satellite imagery to identify anomalous sea level rise at ports on the east coast of the US.
@@
The ensuing paper presents the details of the challenge. In Section 2, we outline the motivation for the challenge and the scientific problem of anomaly detection. We follow this with more detailed descriptions of the datasets and evaluation metrics in Section 3. In Section 4, we present the various datasets comprising the greater challenge. Finally, we discuss how the challenge was made FAIR in Section 5, and we end with the conclusions in Section 6.
2 Motivation
Machine Learning (ML) challenges often consist of well-defined tasks aimed at understanding a general feature of data through either the prediction of a trend in ML regression or the discrimination of different labeled categories through ML classification [32, 33]. Recently, generative AI has added further tasks, such as text or image generation [34, 35]. These methods that are developed for the aforementioned tasks all involve a known objective. The known objective enables a well-defined training strategy. In classifications and regression, this involves the use of cross-entropy and mean squared error minimization through the use of truth-labeled data, respectively. In generation tasks, such as next token prediction, the strategy involves predicting the next or masked token in a series of data. Anomaly detection, in contrast, requires a different training and application strategy since anomalies consist of unlabeled, unknown outliers within the data. Here, we aim to predict our certainty that a given data point aligns with conventional scientific data patterns. Predicting certainty is a much more difficult task since it involves codifying our certainty about the data and what constitutes an anomaly.
A large variety of approaches exist to explore anomaly detection. The simplest of these approaches involves observation of points to see if they deviate from a trend. Traditionally, ML approaches to anomaly detection have been driven by unsupervised learning [21, 36, 17, 18, 37]. In unsupervised learning, one attempts to take a dataset and compress it to a latent space aimed at capturing the salient features that describe the dataset. This is often done by trying to reproduce the dataset or a part of the dataset, similar to next token prediction. Anomaly detection is then performed by looking to see if the projected sample within the latent space falls within the known regions of the compressed latent space (or not, thus implying an anomaly). The ambiguity of anomaly detection arises with what is labeled as an anomaly within the latent space. Anomalies can include noise artifacts/variations, often referred to as glitches, in the data alongside scientifically meaningful anomalies. A strategy to differentiate such glitches from meaningful anomalies is critical to ensure that algorithms are capable of detecting anomalies.
@@ 2 Motivation
The anomaly detection challenges proposed in this paper build on the tradition of anomaly detection by providing unique, challenging elements that differentiate these challenges from previous challenges. With the genomics problem, we utilize subtle features of images as a way to replace the need for tedious DNA analysis to differentiate various species and subspecies. With the gravitational-wave challenge, we aim to codify all allowed laws of astrophysics within a time series to enhance the identification of unknown anomalies. With the climate science problem, we replace dedicated sensor data with large-scale, coarse satellite imagery and utilize this to find deviations from the norm. More conventional anomaly detection aims to just identify outliers; with each challenge here, we add a scientific element that challenges the ability of conventional approaches. The challenge of codifying each scientific element within machine learning adds an additional layer of complexity and highlights the uniqueness of this work.
@@
For each of the components of this challenge, we aimed to construct anomalies in a controlled manner by utilizing scientific knowledge or data that falls outside of the scope of the challenge. With the genomics problem, this involved using DNA data. For the climate science problem, we utilized local water level sensors. For the astrophysics problem, we used our knowledge of astrophysical phenomena to simulate hypothetical anomalous signals that have not been observed within the data. With each of these challenges, our use of hidden labels, sensors and simulations was critical to the dataset construction. However, in real-world scenarios this additional information is not present nor easily obtained,. As a result, devising a strategy that can rely on the limited information—even if not perfect—would lead to impactful scientific results applicable to a large amount of unexplored data. Moreover, through the use of this additional information, we can develop well-defined scoring metrics that make it possible to determine a winner and ensure FAIR principles are enforced throughout the challenge.

### v2

@@ Abstract
Scientific discoveries are often made by finding a pattern or object that was not predicted by the known rules of science. Oftentimes, these anomalous events or objects that do not conform to the norms are an indication that the rules of science governing the data are incomplete, and something new needs to be present to explain these unexpected outliers. Finding anomalies can be confounding since it requires codifying a complete knowledge of the known scientific behaviors and then projecting these known behaviors on the data to look for deviations. When utilizing machine learning, this presents a particular challenge since we require that the model not only understands scientific data perfectly but also recognizes when the data is inconsistent and outside the scope of its trained behavior. In this paper, we present three datasets aimed at developing machine learning-based anomaly detection for disparate scientific domains covering astrophysics, genomics, and polar science. We provide a scheme to make machine learning challenges around the three datasets Findable, Accessible, Interoperable, and Reusable (FAIR). Furthermore, we present an approach that generalizes to future machine learning challenges, enabling the possibility of large, more compute-intensive challenges that can ultimately lead to scientific discovery.
@@ 1 Introduction
Anomaly detection, the process of finding an outlier in a distribution, has long been a critical element of scientific discovery. New anomalies often trigger the question of how to construct a broadly inclusive model allowing for a deeper understanding of the data and, ultimately, a better scientific model. In some cases, this has happened with the observation of a new event or process. Other times, discoveries have been made through aggregating data to observe unexplainable trends [1, 2, 3, 4].
@@
In this paper, we present a series of machine learning (ML) challenges targeting anomaly detection in different scientific domains. We define this as the ability to identify events, objects, or phenomena that are rare and do not match ordinary detector noise or otherwise known signals or phenomena. Anomaly detection is thus searching for the “unknown,” including “known-unknowns” (hypothesized and anticipated but otherwise poorly modeled or quantified) and “unknown-unknowns” (not accommodated by prior knowledge in any given domain that—if detected—push the frontier in understanding the theory and phenomenology of the underlying observations and measurements). This could be the identification of poorly modeled phenomena, such as a core-collapse supernova signal in a gravitational-wave detector[5, 6, 7, 8, 9], unexplained sea ice retreat [10], extreme snow melts in the Arctic [11], or a new species of animal [12, 13, 14]. The challenges presented here aim to bring awareness to the difficulty of using ML for these types of problems: the options of well-labeled datasets or even the possibility of training through a systematic procedure, such as token masking, are often insufficient to answer this type of question. Despite that, there is tremendous potential for algorithms focused on anomaly detection to enable scientific discovery.
@@
The process of using ML for anomaly detection, sometimes referred to as outlier detection, has a long history [15, 16, 17, 18]. However, success has been limited due to the ambiguity of quantifying scientifically meaningful deviation [19, 20]. Nevertheless, several approaches have been tried [20, 18, 21], including the use of unsupervised algorithms [22], which aim to build a mapping for what is regular through an embedding in a latent space, semi-supervised algorithms, which aim to use partially labeled data to construct representation space characteristic to the problem at hand [23], and weakly-supervised algorithms, which aim to play regions of data off of each other to find the largest incongruous regions of the data[24].
@@
The ML challenge project brings together communities involved as part of the NSF Harnessing the Data Revolution (HDR) program, dedicated to connecting domain science with AI to enable groundbreaking scientific discoveries. All five HDR institutes were engaged in developing datasets with a tractable problem defined across their domains, each targeting a single anomaly detection problem and dataset. Ultimately, we narrowed the challenge to three datasets to ensure that they were sufficiently aligned that the same algorithmic toolkit was applicable to each of the challenges. Lastly, the most critical component of this challenge was the ability to hide pre-identified anomalies in such a way that there was no ambiguity in the anomaly definition in the final dataset. The challenge is to see if the proponents can find these pre-hidden anomalies with the smallest possible false positive rate.
The three challenges cover a wide variety of data types and problems, ranging from time series data, to photographic images, and satellite imagery over time, as described below.
@@
  • Anomalous Gravitational-Wave Transients: Short-duration (O(1s)) astrophysical signals present in public data from the two LIGO gravitational-wave detectors. The data that is received from each of the LIGO interferometers is a gravitational-wave strain from a signal, indicating a fractional stretching and squeezing of space-time on the order of $10^{-22}$ the measured length scale. The challenge is to find a signal that appears in both detectors and has broad features representative of an astrophysical signature. Such a signature may correspond, for example, to a core-collapse supernova, a white noise burst of unknown origin, or a black hole merger beyond our nominal modeling and expectation from general relativity.
@@
  • Butterfly Hybrids: Images of butterflies from two mimetic species, each with highly varied subspecies [25]. It includes images of the various subspecies, as well as different “hybrids” among each species’ subspecies. The challenge is to find the anomalous butterfly hybrids without confusing them for subspecies of the other species.
@@
  • Anomalous Sea Level Rise: Satellite imagery [26] depicting observed sea levels over time. The challenge is to use this imagery to identify anomalous sea level rise at ports on the east coast of the US.
A diagram of the challenge flow is provided in Figure 1. For the challenge, teams must develop an algorithm that processes the data and outputs an anomaly score. The scoring is then determined by seeing if the algorithm is capable of singling out the anomalous features of the dataset. Each challenge has three core data subsets:
  • Training Dataset: A publicly released subset consisting of the background (non-anomalous) cases and sample anomalous signals to enable the model construction. These signals do not encompass all variations of the challenge dataset, instead providing examples of what might be present. Moreover, the background dataset need not have the same proportional composition as the challenge dataset.
  • Testing Dataset: During the competition, this subset is used for benchmarking algorithms and presenting a leaderboard on the competition platform. It largely consists of background non-anomalous events with some hidden anomalies (known-unknowns identifiable through additional information not available to participants). This dataset is not released publicly; however, participants may run their algorithm on it through the collaboration platform.
  • Challenge Dataset: This subset is used to determine the winner. It is constructed in the same way as the testing dataset, but is not used before the final scoring. Thus, no participating team can overly tune the algorithm on the dataset.
Figure 1: Overview of the flow of the ML challenge, and the datasets used for training. All preparation was explicitly made following FAIR principles, with the curation and data cleaning workflows stored on GitHub. The Codabench platform and NERSC computing cluster are used for scoring and ranking of submissions.
A driving goal of this work is to codify a list of best practices that ensure workflows, no matter the scientific domain, are easy to follow and reproduce. Through this work, we endeavor to establish an exemplar for future ML challenges, preserving the level of Findability, Accessibilty, Interoperability, and Reusability (FAIR) present in this challenge in future challenges. Moreover, by ensuring a high level of FAIRness we aim to use these datasets for future benchmarking[19, 27, 28].
@@
The ensuing paper presents the details of the challenge. In Section 2, we outline the motivation for the challenge and the scientific problem of anomaly detection. Subsection 2.1 highlights how the challenge was made FAIR. We then present the various datasets comprising the greater challenge in Section 3, followed by a detailed description of the evaluation metric in Section 4, and we end with the conclusions.
2 Scientific and FAIR Motivation
@@
The anomaly detection challenges presented in this paper build on the traditional anomaly detection by providing unique, scientific elements that distinguish them from previous challenges. With the genomics problem, we utilize subtle features of images to replace the need for tedious DNA analysis to differentiate various species and subspecies. With the gravitational-wave challenge, we aim to codify all allowed laws of astrophysics within a time series to enhance the identification of unknown anomalies. With the climate science problem, we replace dedicated sensor data with large-scale, coarse satellite imagery and utilize this to find deviations from the norm. More conventional anomaly detection simply targets outliers; with each challenge here, we add a scientific element that pushes the boundaries of conventional approaches. The challenge of codifying scientific elements within ML adds an additional layer of complexity and highlights the uniqueness of this work.
@@
For each component of this challenge, we constructed anomalies in a controlled manner, utilizing scientific knowledge or data not provided in the challenge. With the genomics problem, this involved using DNA. For the climate science problem, we utilized local water level sensors. For the astrophysics problem, knowledge of astrophysical phenomena is used to simulate hypothetical anomalous signals not observed within the data. With each of these challenges, our use of hidden labels, sensors and simulations is critical to the dataset construction. However, in real-world scenarios this additional information is not present, nor easily obtained. As a result, devising a strategy that can rely on the limited information—even if imperfect—would lead to impactful scientific results applicable to a large amount of unexplored data.